= 100 + 36Q-4a-and C(Q-80 + 12Q. (Note: MBQ)-36-8Q and MCQG) -12) Instructions: Use a negative sign () where appropriate. a. Write out the equation for the net benefits. b, what are the net benefits when Q-1? Q-5? Net benefits when Q 1 Net benefits when Q 5 C. Write out the equation for the marginal net benefits. d, what are the marginal net benefits when Q#1? Q-57 Marginal net benefits when Q 1: Marginal net benefits when Q-5: e. What level of Q maximizes net benefits? f At the value of Q that maximizes net benefits, what is the value of marginal net benefits?
Answer #1

Answer:

1)

B=100+36Q-4Q^2

C=80+12Q

Net Benefit N=B-C=100+36Q-4Q^2-(80+12Q)

N=20+24Q-4Q^2

2)

when Q=1

N=20+24*1-4*1^2=$40

when Q=5

N=20+24*5-4*5^2=$40

3)

MB=36-8Q

MC=12

MNB=MB-MC=36-8Q-12

MNB=24-8Q

4)

MNB when Q=1

MNB=24-8*1=$16

when Q=5

MNB=24-8*5=-$16

5)

For maximization of net benefit

MB=MC

36-8Q=12

Q=3 units

6)

for Q=3

Marginal Net benefit

MNB=24-8*3=0

Net Benefit

N=20+24Q-4Q^2=20+24*3-4*3^2=$56